### v1

@@ 6 Experiments
Once a solution to the Nonlinear Modified PageRank problem is available, graph partitioning can take place as outlined in Section 2.2. Results pertaining to the local cluster quality on both synthetic and real-world inspired graphs are now reported.
@@ 6.1 Set-up
An implementation consisting of the Nonlinear Modified PageRank problem numerical solution and sweep-cut was made in MATLAB R2022b. immoptibox [25] was adopted for the featured Levenberg-Marquardt method. However, adjustments to the routine were required in order to accommodate the full rank Jacobian variant. Dijkstra’s algorithm [8] was called on for determining the furthest distance vertex from the starting vertex $s\in\mathcal{V}$. The source code can be found at https://github.com/DmsPas/Nonlinear_modified_PageRank.
For the problem solution, $\beta=0.01$ except where stated otherwise. The value at the vertex judged to be the furthest was $10^{-12}$. $\zeta$ was either $10^{-6}$ for graphs with $n\geq 10^{4}$, or $10^{-11}$. Termination criteria in the form of a gradient (maximum) norm and a relative change in the potential solution were set to $10^{-7}$.
It is the local cluster with the smallest conductance at any value of $p$ that is of interest. Accordingly, the Nonlinear Modified PageRank problem defined on a graph $\mathcal{G}$ was tackled for $p=1.95$, $1.9$, $1.8$, $1.7$, $1.6$, $1.5$, and $1.45$ sequentially (these values were selected based on experiment). Each solution served as the initial iterate $x_{0}$ for the next optimisation problem with the exception obviously being $p=1.95$. The initial iterate for $p=1.95$ was the minimal least squares solution to the Nonlinear Modified PageRank problem at $p=2$. $\lambda_{0}$ was taken to be the product of $10^{-3}$ and the largest entry in the main diagonal of $\widetilde{J}(x_{0})^{\top}\widetilde{J}(x_{0})$.
Local clusters based on the Nonlinear Modified PageRank, which is abbreviated to NPR, problem solution were compared to those from sweep-cuts of solutions resulting from the following:
  • Approximate Personalised PageRank (APPR) [12]. An efficient algorithm for computing accurate approximations of the PageRank problem (5). Even though the algorithm was designed with web graphs in mind, usage here was primarily for validation purposes.
@@
Source code for the competing methodologies is in the public domain. Suggested operating parameter settings were unchanged. For all graphs, a starting vertex was selected at random, unless otherwise specified, and then the conductance with respect to $p$ was investigated. This was repeated a total of 50 times for every synthetic graph, and 10 times for the real-world inspired graphs.
Vertices in every graph, bar none, were assigned a binary ground truth label, making it possible to easily identify those that constitute a local cluster. The labels supported quality assessment of discovered clusters through a measure called (balanced) $Fscore\in[0,1]$, which is defined as
@@
where $precision=TP/(TP+FP)$ and $recall=TP/(TP+FN)$. $TP$ represents the number of true positives, i.e., vertices correctly classified in the local cluster. $FP$ is the number of false positives, i.e., vertices incorrectly classified in the local cluster. $FN$ stands for the number of false negatives, i.e., vertices that should have been part of the local cluster but were not. A value of $Fscore=1$ indicates perfect recovery of a local cluster. Mean and standard deviation of both the conductance and related $Fscore$ were recorded as standard, that is, where relevant.
[…]
0.10.10.10.10.20.20.20.20.30.30.30.30.40.40.40.40.20.20.20.20.40.40.40.4Mixing parameter ($\gamma$)Conductance ($\Phi$)
0.10.10.10.10.20.20.20.20.30.30.30.30.40.40.40.4000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81111Mixing parameter ($\gamma$)$Fscore$ (corresponding to $\Phi$)
Figure 1: (a) Conductance and the related (b) $Fscore$ of local clusters for an increasing number of inter-community edges in LFR graphs.
@@ 6.2.1 Communities benchmark
The Lancichinetti–Fortunato–Radicchi (LFR) model [20] produces non-weighted graphs with communities. Vertex degree and community size distributions in the graphs correspond to power laws. A mixing parameter $\gamma$ in the model is responsible for inter-community edges. The greater the value of $\gamma$, the more difficult it is to distinguish independent communities.
For the purpose of graph generation, the number of vertices was 1,000. The average vertex degree was 10 and the maximum was 50. With regard to the community size, the minimum was 20 vertices and the maximum was 100. The power law exponents for the vertex degree distribution and community size distribution were two and one, respectively. Values of $\gamma$ were allowed to range from 0.1 to 0.4.
@@
Figure 2: Levenberg-Marquardt method for the NPR problem defined on an LFR graph with $\gamma=0.3$ at (a) $p=1.95$, (b) $p=1.8$, and (c) $p=1.6$.
Plots in Figure 2 demonstrate the performance of the Levenberg-Marquardt method for solution of the NPR problem defined on a graph with $\gamma=0.3$ at $p=1.95,1.8$ and $1.6$. It can be seen that the number of iterations to achieve convergence increases slightly as $p$ goes to $1.6$. Note the decrease in mean conductance from $0.404$ to $0.359$, and the increase in mean $Fscore$ from $0.78$ to $0.83$.
[…]
0010101010202020203030303040404040000.020.020.020.020.040.040.040.04Number of groupingsConductance ($\Phi$)
[…]
Figure 3: (a) Example of eight groupings. (b) Local cluster based on the NPR problem solution. The starting vertex is shown as a yellow diamond, and the edges comprising the boundary are in red. (c) Conductance and the related (d) $Fscore$ of local clusters for an increasing number of groupings.
@@ 6.2.2 Gaussian datasets
Groupings comprised of 400 points in $\mathbb{R}^{2}$ randomly sampled (in each dimension) from a normal distribution with a variance of 0.055 were generated. Their centres were located equidistantly on a square grid. Connectivity between points was established by a nearest neighbour search set to 10 points, i.e., the 10 closest (distance-wise) points to every point were considered connected to that point.
Graphs representing 2, 5, 8, 13, 18, 25, 32, and 41 groupings were constructed. While points are the geometric manifestation or realisation of vertices, edges correspond to the connections between points. The number of edges in each graph was 4,902, 12,153, 19,319, 31,204, 43,013, 59,644, 76,044, and 97,112. Let points be denoted by $y$. For every graph $\mathcal{G}$ and all $\{u,v\}\in\mathcal{E}$, the weight was given by
@@
(c) $p=1.60$
Figure 4: Levenberg-Marquardt method for the NPR problem defined on a graph with eight groupings at (a) $p=1.95$, (b) $p=1.8$, and (c) $p=1.6$.
Plots in Figure 4 demonstrate the performance of the Levenberg-Marquardt method for solution of the NPR problem defined on a graph accounting for eight groupings at $p=1.95,1.8$, and $1.6$. As opposed to before, the number of iterations to achieve convergence actually decreases a little as $p$ goes to $1.6$. Note the decrease in mean conductance from $0.014$ to $0.012$, and the increase in mean $Fscore$ from $0.82$ to $0.87$.
@@ 6.3 Real-word inspired graphs
6.3.1 Image databases
There is quite the number of image databases that have been developed in order to facilitate research in pattern recognition and machine learning. The MNIST database [7] of handwritten digits from zero to nine has been used extensively. A test set of 10,000 $28\times 28$ grayscale images is provided as part of the database. Similarly, the USPS database [14] contains 11,000 handwritten digit $16\times 16$ grayscale images that were scanned from envelopes in a working post office (by the U.S. Postal Service). Fashion-MNIST [37] is a database that is intended to pose a more significant challenge than the handwritten digits in MNIST. The database offers a test set of 10,000 $28\times 28$ grayscale images from (online retailer) Zalando’s website. Each image is categorised into one of ten classes: t-shirt/top, trouser, pullover, dress, coat, sandals, shirt, sneaker, bag, and ankle boots.
Table 1: Mean conductance and the related mean $Fscore$ of local clusters representing subsets of images in the following databases: MNIST, Fashion-MNIST, and USPS.
@@ 6.3.1 Image databases
$\Phi$ $Fscore$ $\Phi$ $Fscore$ $\Phi$ $Fscore$
@@
APPR $0.282\scriptstyle{\pm 5\cdot 10^{-2}}$ $0.317\scriptstyle{\pm 1\cdot 10^{-1}}$ $0.234\scriptstyle{\pm 4\cdot 10^{-2}}$ $0.279\scriptstyle{\pm 2\cdot 10^{-1}}$ $\mathbf{0.071}\scriptstyle{\pm 1\cdot 10^{-2}}$ $0.211\scriptstyle{\pm 2\cdot 10^{-2}}$
@@
$p$-DIFF $0.111\scriptstyle{\pm 3\cdot 10^{-2}}$ $0.336\scriptstyle{\pm 1\cdot 10^{-1}}$ $0.043\scriptstyle{\pm 1\cdot 10^{-2}}$ $0.317\scriptstyle{\pm 9\cdot 10^{-2}}$ $0.116\scriptstyle{\pm 2\cdot 10^{-2}}$ $0.312\scriptstyle{\pm 2\cdot 10^{-2}}$
Images can be treated as points in Euclidean space of dimension equal to their resolution (i.e., the total number of pixels). Subsequently, graphs were constructed for the aforementioned databases as was done previously. Table 1 presents the mean conductance and related mean $Fscore$ associated with local clusters. Those based on the NPR problem solution have the highest mean $Fscore$ and the lowest conductance in all but one case, which is the USPS database.
@@
Figure 5: Local clusters based on the NPR problem solution in black and blue, and their ground truth equivalent with the same number of vertices in black and red. The number of vertices belonging to a local cluster is (a) 24, (b) 38, (c) 151, (d) 12, (e) 42, (f) 165, (g) 58, (h) 33, and (i) 76. Yellow diamonds represent Constantinopolis (a,d,g), Londinium (b,e,h), and Roma (c,f,i). Clusters are related to the distance (a,b,c), duration (d,e,f), and financial cost (g,h,i) between sites.
6.3.2 Roman world communication network circa 200 AD
@@
Three graphs were constructed, each with 677 vertices and 1104 edges representing sites and travel routes, respectively. They differed only in weight function definition. For every graph $\mathcal{G}$ and all $\{u,v\}\in\mathcal{E}$, the weight was given by
@@
where $\chi$ is the distance, duration, or financial cost between the sites represented by vertices $u$ and $v$. The raw data from ORBIS can have a value for one direction between vertices $u$ and $v$, and another for the opposition direction. An average was, therefore, taken between two values in all cases. $\iota$ is the mean value of the distance, duration, or travel cost between $u$ and adjacent vertices.
@@
Figure 5 visualises the local clusters and their ground truth counterparts. Starting vertices are illustrated by yellow diamonds. Sites coloured black are those that are part of both clusters. While blue sites are those that were not part of the ground truth cluster, the red sites are those that were only part of the ground truth cluster. Other than for two cases, the clusters based on the NPR problem solution were in very good to excellent agreement with the ground truth clusters. Even for those two cases, performance was still respectable with an $Fscore=0.709$ and $Fscore=0.672$. An $Fscore=1$ was registered for the graph with the travel duration edge weight and the starting vertex of Constantinopolis.
@@ Acknowledgments
Both authors acknowledge the scientific support and HPC resources provided by the Erlangen National High Performance Computing Center (NHR@FAU) of the Friedrich-Alexander-Universität Erlangen-Nürnberg (FAU). NHR funding is provided by federal and Bavarian state authorities. NHR@FAU hardware is partially funded by the German Research Foundation (grant 440719683). A great debt of gratitude is owed to Prof. Horia Cornean for his review of the manuscript. His advice and support significantly enhanced the quality of the article.

### v2

@@ 6 Experiments
Once a solution to the Nonlinear PageRank problem is available, graph partitioning can take place as outlined in Section 2.2. Results pertaining to the local cluster quality on both synthetic and real-world inspired graphs are now reported.
@@ 6.1 Set-up
An implementation consisting of the Nonlinear PageRank problem numerical solution and sweep-cut was made in MATLAB R2024b. immoptibox [30] was adopted for the Levenberg-Marquardt method with adjustments to accommodate the full rank Jacobian variant. Dijkstra’s algorithm [31] was called on for determining the furthest distance vertex from the starting vertex. The source code can be found at https://github.com/DmsPas/Nonlinear_modified_PageRank.
For the problem solution, $\beta=0.01$ unless specified differently. The value at the vertex judged to be the furthest was $10^{-12}$. $\zeta$ was either $10^{-11}$ for graphs with fewer than $10^{4}$ vertices or $10^{-6}$ otherwise. Termination criteria in the form of a gradient (maximum) norm and a relative change in the potential solution were set to $10^{-7}$.
In order to determine the local cluster that has the smallest conductance with respect to $p$, the Nonlinear PageRank problem defined on a graph $\mathcal{G}$ was tackled for $p=1.95$, $1.9$, $1.8$, $1.7$, $1.6$, $1.5$, and $1.45$ sequentially (these values were selected based on experiment). Each solution served as the initial iterate $x_{0}$ for the next optimisation problem with the exception of $p=1.95$, which was the minimal least squares solution to the Nonlinear PageRank problem at $p=2$. $\lambda_{0}$ was taken to be the product of $10^{-3}$ and the largest entry in the main diagonal of $\widetilde{J}(x_{0})^{\top}\widetilde{J}(x_{0})$.
Local clusters based on the Nonlinear PageRank (hereafter abbreviated as NPR) problem solution were compared to those from sweep-cuts of solutions resulting from the following:
  • Approximate Personalised PageRank (APPR) [32]. An efficient algorithm for computing accurate approximations of the PageRank problem (2).
@@
Source code for the competing methodologies is in the public domain. Operating parameter settings were maintained in the experiments. For all graphs, a starting vertex was selected at random, unless otherwise specified, and the conductance of the local cluster formed at each $p$ was ascertained. This was repeated a total of 50 times for every synthetic graph, and 10 times for the real-world inspired graphs.
Vertices in every graph were assigned a binary ground truth label to identify those that constitute a local cluster. The labels supported quality assessment of the discovered clusters through a measure called (balanced) $Fscore\in[0,1]$, which is defined as
@@
where $precision=TP/(TP+FP)$ and $recall=TP/(TP+FN)$. TPTP represents the number of true positives, i.e., vertices correctly classified in the local cluster. $FP$ is the number of false positives, i.e., vertices incorrectly classified in the local cluster. $FN$ stands for the number of false negatives, i.e., vertices that should have been part of the local cluster but were not. A value of $Fscore=1$ indicates perfect recovery of a local cluster. Mean and standard deviation of both the conductance and related $Fscore$ were recorded as standard, that is, where relevant.
NPRAPPRNPDpp-DIFF
0.10.10.20.20.30.30.40.40.20.20.40.4Mixing parameter ($\gamma$)Conductance ($\Phi$)
0.10.10.20.20.30.30.40.400.20.20.40.40.60.60.80.811Mixing parameter ($\gamma$)FscoreFscore (corresponding to $\Phi$)
Figure 1: (a) Conductance and the related (b) FscoreFscore of local clusters for an increasing number of inter-community edges in LFR graphs.
@@ 6.2.1 Communities benchmark
The Lancichinetti–Fortunato–Radicchi (LFR) model [34] produces graphs with known communities. Vertex degree and community size distributions in the graphs follow power laws. A mixing parameter $\gamma$ is responsible for inter-community edges. The greater the value of $\gamma$, the more difficult it is to distinguish independent communities.
For the purpose of (non-weighted) graph generation, the number of vertices was 1,000 with an average vertex degree of 10 and a maximum of 50. Community sizes ranged from 20 to 100 vertices. The power law exponents for the vertex degree distribution and community size distribution were set to 2 and 1, respectively. Values of $\gamma$ were varied from 0.1 to 0.4.
@@
Figure 2: Levenberg-Marquardt method for the NPR problem defined on an LFR graph with $\gamma=0.3$ at (a) $p=1.95$, (b) $p=1.8$ and (c) $p=1.6$.
Plots in Figure 2 illustrate the convergence history of the Levenberg-Marquardt method for solution of the NPR problem defined on a graph with $\gamma=0.3$ at $p=1.95,1.8$ and $1.6$. It can be seen that the number of iterations to achieve convergence increases slightly as $p$ goes to $1.6$. Note the decrease in mean conductance from $0.404$ to $0.359$ and the increase in mean $Fscore$ from $0.78$ to $0.83$.
NPRAPPRNPDpp-DIFF
02020404000.020.020.040.04Number of groupingsConductance ($\Phi$)
02020404000.20.20.40.40.60.60.80.811Number of groupingsFscoreFscore (corresponding to $\Phi$)
Figure 3: (a) Example of eight groupings. (b) Local cluster based on the NPR problem solution. The starting vertex is shown as a yellow diamond and the edges comprising the boundary are in red. (c) Conductance and the related (d) FscoreFscore of local clusters for an increasing number of (400 point) groupings.
@@ 6.2.2 Gaussian datasets
Groups of 400 points in $\mathbb{R}^{2}$ randomly sampled (in each dimension) from a normal distribution with a variance of 0.055 were generated and their centres located equidistantly on a square grid. Connectivity between points was established by a nearest neighbour search set to 10 points, i.e., the 10 closest (distance-wise) points to every point were considered connected to that point.
Graphs representing 2, 5, 8, 13, 18, 25, 32 and 41 groupings were constructed. While points are the geometric manifestation or realisation of vertices, edges correspond to the connections between points. The number of edges in each graph was 4,902, 12,153, 19,319, 31,204, 43,013, 59,644, 76,044 and 97,112. Let points be denoted by $y$. For every graph $\mathcal{G}=(\mathcal{V},\mathcal{E},w)$ and all $\{u,v\}\in\mathcal{E}$, the assigned weight was given by
@@
(c) $p=1.6$
Figure 4: Levenberg-Marquardt method for the NPR problem defined on a graph with eight groupings at (a) $p=1.95$, (b) $p=1.8$ and (c) $p=1.6$.
Plots in Figure 4 demonstrate the performance of the Levenberg-Marquardt method for solution of the NPR problem defined on a graph accounting for eight groupings at $p=1.95,1.8$ and $1.6$. As opposed to before, the number of iterations to achieve convergence actually decreases a little as $p$ goes to $1.6$. Note the decrease in mean conductance from $0.014$ to $0.012$ and the increase in mean $Fscore$ from $0.82$ to $0.87$.
@@ 6.3 Real-word inspired graphs
6.3.1 Image classification
There is quite the number of image databases that have been developed in order to facilitate research in pattern recognition and machine learning. The MNIST database [35] of handwritten digits from zero to nine has been used extensively. A test set of 10,000 $28\times 28$ grayscale images is provided as part of the database. Similarly, the USPS database [36] contains 11,000 handwritten digit $16\times 16$ grayscale images that were scanned from envelopes in a working post office (by the U.S. Postal Service). Fashion-MNIST [37] offers a test set of 10,000 $28\times 28$ grayscale images of clothing items from (online retailer) Zalando’s website. Each image is categorised into one of ten classes: t-shirt/top, trouser, pullover, dress, coat, sandals, shirt, sneaker, bag and ankle boots.
Table 1: Mean conductance and the related mean FscoreFscore of local clusters representing subsets of images in the following databases: MNIST, Fashion-MNIST and USPS.
@@ 6.3.1 Image classification
$\Phi$ FscoreFscore $\Phi$ FscoreFscore $\Phi$ FscoreFscore
@@
[…]
@@
$p$-DIFF $0.111\scriptstyle{\pm 3\cdot 10^{-2}}$ $336\scriptstyle{\pm 1\cdot 10^{-1}} 0.0$$043\scriptstyle{\pm 1\cdot 10^{-2}}$ $0.317\scriptstyle{\pm 9\cdot 10^{-2}}$ $0.116\scriptstyle{\pm 2\cdot 10^{-2}}$ $0.312\scriptstyle{\pm 2\cdot 10^{-2}}$
Images can be treated as points in Euclidean space of dimension equal to their resolution (i.e., the total number of pixels). As such, graphs were constructed for the aforementioned databases as was done previously. Table 1 presents the mean conductance and related mean $Fscore$ associated with local clusters. Those based on the NPR problem solution have the highest mean $Fscore$ and the lowest conductance in all but one case, which is the USPS database.
@@
Figure 5: Local clusters based on the NPR problem solution in black and blue, and their ground truth equivalent with the same number of vertices in black and red. The number of vertices belonging to a local cluster is (a) 24, (b) 38, (c) 151, (d) 12, (e) 42, (f) 165, (g) 58, (h) 33 and (i) 76. Yellow diamonds represent Constantinopolis (a,d,g), Londinium (b,e,h) and Roma (c,f,i). Clusters are related to the distance (a,b,c), duration (d,e,f) and financial cost (g,h,i) between sites.
6.3.2 Roman world transport network circa 200 AD
@@
Three graphs were constructed, each with 677 vertices and 1104 edges representing sites and travel routes, respectively. They differed only in weight function definition. For every graph $\mathcal{G}=(\mathcal{V},\mathcal{E},w)$ and all $\{u,v\}\in\mathcal{E}$, the assigned weight was given by
@@
where $\chi(u,v)$ is the distance, duration, or financial cost between the sites represented by vertices $u$ and $v$. The raw data from ORBIS can have a value for one direction between vertices $u$ and $v$, and another for the opposition direction. An average was, therefore, taken between two values in all cases. $\iota$ is the mean value of the distance, duration, or financial cost between $u$ and adjacent vertices.
@@
Figure 5 visualises the local clusters and their supposed ground truth counterparts. Starting vertices are illustrated by yellow diamonds. Sites coloured black are those that are part of both clusters. While blue sites are those that were not part of the ground truth cluster, the red sites are those that were only part of the ground truth cluster. Other than for two cases, the clusters based on the NPR problem solution were in very good to excellent agreement with the ground truth clusters. Even for those two cases, performance was still respectable with an $Fscore=0.709$ and $Fscore=0.672$. An $Fscore=1$ was registered for the graph with the travel duration edge weight and the starting vertex of Constantinopolis.
@@ Acknowledgments
The authors gratefully acknowledge the scientific support and HPC resources provided by the Erlangen National High Performance Computing Center (NHR@FAU) of the Friedrich-Alexander-Universität Erlangen-Nürnberg (FAU) under the NHR project 80227. NHR funding is provided by federal and Bavarian state authorities. NHR@FAU hardware is partially funded by the German Research Foundation (DFG) – 440719683. A great debt of gratitude is owed to Prof. Horia Cornean for his review of the manuscript. His advice and support significantly enhanced the quality of the article.